Caspian Sea Drinks is considering buying the J-Mix 2000. It will allow them to make and sell more product. The machine cost $1.52 million and create incremental cash flows of $523,661.00 each year for the next five years. The cost of capital is 8.93%. What is the internal rate of return for the J-Mix 2000?

Derek plans to buy a $30,262.00 car. The dealership offers zero percent financing for 57.00 months with the first payment due at signing (today). Derek would be willing to pay for the car in full today if the dealership offers him $____ cash back. He can borrow money from his bank at an interest rate of 5.02%.
Answer #1

1: IRR = 21.37%

2: Cash back = 30262-23853.72 = 6408.28

PV of the purchase = FV/(1+r)^n

= 30262/(1+5.02%/12)^57

= 23853.72

Workings

AutoSaveCff) Book1 Excel Sign in Comm View O Tell me what you want to do File Insert Formulas Review Help Share Home Draw Pag